A bottle of acetone (fingernail polish remover) contains 43.5 mL of acetone with a mass of 55.0 g.
Express the density in grams per milliliter to three significant figures.
1
Expert's answer
2019-09-05T06:17:01-0400

Solution. p= m/V= 1.264 (g/ml)

Answer. 1.26 g/ml